### v1

# Hierarchical LLMs In-the-loop Optimization for Real-time Multi-Robot Target Tracking under Unknown Hazards

Abstract
In this paper, we propose a hierarchical Large Language Models (LLMs) in-the-loop optimization framework for real-time multi-robot task allocation and target tracking in an unknown hazardous environment subject to sensing and communication attacks. We formulate multi-robot coordination for tracking tasks as a bi-level optimization problem, with LLMs to reason about potential hazards in the environment and the status of the robot team and modify both the inner and outer levels of the optimization. The inner LLM adjusts parameters to prioritize various objectives, including performance, safety, and energy efficiency, while the outer LLM handles online variable completion for team reconfiguration. This hierarchical approach enables real-time adjustments to the robots’ behavior. Additionally, a human supervisor can offer broad guidance and assessments to address unexpected dangers, model mismatches, and performance issues arising from local minima. We validate our proposed framework in both simulation and real-world experiments with comprehensive evaluations, which provide the potential for safe LLM integration for multi-robot problems.
I Introduction
Multi-robot coordination and planning have been extensively employed in different applications like exploration, search and rescue, and target tracking [1, 2, 3, 4, 5]. Coordinating a team of robots in unknown and adversarial environments becomes significantly challenging due to the increased complexity and unpredictability of conditions. Moreover, sensor and communication failures, potentially caused by adversaries, can lead to critical safety issues, including the inability of robots to localize, observe their surroundings, and maintain team connectivity. While existing approaches have focused on designing algorithms to handle modeled adversarial tasks and environments [6, 7, 8, 9], the mismatch between these models and the unexpected dangers encountered in real-world scenarios makes adaptation difficult.
Figure 1: Hardware experiment of multi-drone target tracking with human input. Two drones are tracking two ground robots within an environment that features two sensing danger zones (in red) and one communication danger zone (in blue) as potential hazards. A human supervisor offers guidance to adjust the drones’ behavior to mitigate risks.
The recent advancements in Large Language Models (LLMs) have demonstrated great potential for improving task and motion planning for robots [10, 11, 12, 13, 14, 15]. Existing approaches primarily focus on using LLMs for offline applications to provide high-level guidance and rewards for sequential tasks without strict temporal constraints or real-time requirements. For high-level tasks, LLM-based systems typically depend on human experts to offer linguistic or numerical guidance, refining strategies that enable the robot team to navigate complex and uncertain environments effectively [16]. However, mismatches between models and real-world conditions, along with unexpected risks, in multi-robot target tracking can overwhelm human experts due to excessive cognitive demands [17, 18]. Moreover, in many practical applications, it is more common for non-expert human supervisors to watch over the robot team. These supervisors can offer vague natural language suggestions but typically lack the expertise to adjust mathematical models [19, 20, 21]. Therefore, this creates a need for a real-time coordination assistant to continuously monitor the robot team, incorporate human suggestions, and make strategic adaptations to benefit decision-making.
Figure 2: The hierarchical LLMs framework for multi-robot tracking. The robots’ inputs and human inputs are processed by two LLMs: a task LLM and an action LLM. The action LLM operates at a higher frequency for reactive performant adaptation, while the task LLM has a lower frequency to provide strategic guidance and reconfiguration to the system. The optimization solver takes the LLMs’ outputs into executable actions, guiding the multi-robot team in real-time.
To address these challenges, we hierarchically integrate LLMs into multiple feedback loops alongside conventional optimizers, leveraging high-level reasoning and decision-making to combine the strengths of both approaches, as illustrated in Fig. 2. We formulate the task assignment and target tracking problem as a bi-level optimization, utilizing the understanding and reasoning capabilities of LLMs for distinct subproblems. Specifically, the outer loop LLM identifies task allocation variables to decide which target to track with reasoning significance, while the inner loop LLM tunes parameters to prioritize objectives towards tracking performance, safety, and energy efficiency to guide the behaviors of robots. In addition to in-the-loop numerical feedback, we also enable human interactions in the system to address abnormal behaviors, remind unexpected danger avoidance, and provide more high-level supervision. We evaluate the proposed framework with several scenarios and deploy it on our customized hardware platforms to validate its real-time robustness and capabilities.
II Related Works
Compared with traditional multi-robot coordination methods, risk-aware and danger-aware strategies have gained significant attention in recent years for their potential to improve resilience and robustness in real-world applications [22]. Hazards can cause temporary or permanent damage to robot sensors, disrupt communication links between robots, and degrade the performance of task execution. Authors in [6] addressed these dangers in multi-robot target tracking and explored chance-constrained optimization that effectively balances minimizing the risk of failures with maximizing target tracking efficiency. Resiliency and adaptive strategies with recovery from sensor and communication failures were further extended in [23, 7]. To migrate the influence of sensing and communication attacks, the work in [24] employed neural models for the formation control of the robot team. Other methods generated active protections by employing jamming techniques to counter eavesdropping [25]. Additionally, active strategies like visual-based anomaly detection [26, 13] helped ensure safe operations under unforeseen environmental conditions.
However, the environments and adversaries considered in these prior studies are ideal compared to real-world scenarios, which involve challenges such as limited heterogeneous resources, dynamically shifting adversarial zones with changing shapes, and the need for real-time adaptation. With the increasing complexity of tasks and the hazards of the environments, the problem formulation might not always be feasible, and optimal solutions cannot be obtained in allowable time with limited computation resources.
Incorporating LLMs into task allocation and motion planning has shown promising results in areas such as reward design [27], generating subgoals [28, 29], and anomaly detection [13]. While these models can offer valuable strategic input, they are generally ineffective and lack guarantees for managing low-level tasks or direct robot actions when used in isolation, as highlighted in previous research [30, 31, 32]. For more complex scenarios with hard requirements, combining LLMs with conventional optimization techniques can enhance both the feasibility and effectiveness of the solutions. LLMs contribute valuable guidance for iterative refinement and problem formulation adjustments [33, 34], while conventional methods offer stronger guarantees for solution quality and constraint handling. This integration motivates our work to explore the potential for more robust and reliable performance.
III Approach
III-A System Overview
The proposed framework of multi-robot tracking with hierarchical LLMs in-the-loop is shown in Fig. 2. We leverage LLMs in outer and inner feedback loops to iteratively improve the performance for real-time deployment. The outer loop with task LLM evaluates environment changes and task performance, and ensures that the goal assignments are optimal for the robot team. The inner loop runs at a high frequency with direct numerical performance feedback to adjust the priority among tracking performance, energy, and safety toward different types of dangers. With specified variables and parameters, we employ optimizers to generate guaranteed intelligent and safe actions for the robot team. To address high-level environmental and scenario-specific risks, we incorporate real-time human input, such as feedback on robot performance and environmental hazards, to provide more comprehensive and adaptive responses. This ensures that the multi-robot team can navigate safely and efficiently in dynamic and potentially hazardous environments. When the optimized actions are broadcast, the robots execute actions, update their observations, and work together to gather data, detect failures, and provide feedback.
III-B Task Allocation and Planning as a Bi-level Optimization
We utilize the multi-robot multi-target tracking problem as an instance of multi-robot coordination with task and motion planning to showcase the capabilities of our framework. The unknown hazards in the environments are modeled as danger zones with probabilistic attacks on the robot team. The problem is defined as follows.
Figure 3: Demonstration of multi-robot target tracking with sensing and communication danger zones. The robot team consists of five robots tracking seven targets in an environment with four danger zones. In (a), one robot is attacked while inside a sensing danger zone, and in (b), it successfully recovers from the attack. In (c), a communication attack causes another robot to lose its communication link, and in (d), this robot recovers communication and resumes information sharing with the team.
Definition 1.
(Multi-robot Target Tracking with Danger Zones [6]): We consider M robots equipped with sensors to track N moving targets in hazardous environments. These hazards are categorized into two types: sensing danger zones and communication danger zones. The objective of the multi-target tracking problem is to maximize tracking performance while balancing the deteriorating impacts caused by attacks from these danger zones.
This problem requires robots to assign tracking goals toward targets and maximize the tracking performance, which is usually formulated as minimizing the trace of the estimation covariance matrix. A demonstration of the tracking process is shown in Fig. 3. Generally, we formulate this problem of multi-robot coordination in adversarial environments as a constrained multi-objective nonlinear optimization:
$\vspace{-0.4cm}\begin{split}\min_{\mathcal{A},\mathbf{u}}\ \ &\mathcal{J}(\mathcal{A},\mathbf{u})=\{\mathcal{J}_{1}(\mathcal{A},\mathbf{u}),\cdots,\mathcal{J}_{n}(\mathcal{A},\mathbf{u})\},\\ {s.t.}~{}\ &\mathcal{H}(\mathcal{A},\mathbf{u})=\bm{0},\quad\mathcal{G}(\mathcal{A},\mathbf{u})\preceq\bm{0},\end{split}$ (1)
where $\mathcal{J}(\cdot)$ represents a set of objectives, such as tracking performance, energy, and safety considerations. The term $\mathcal{H}(\cdot)$ denotes the general equality constraint functionals, while $\mathcal{G}(\cdot)$ represents the inequality constraint functionals. These constraints encode dynamics-related limitations, feasibility bounds, and other system-specific restrictions. The variables $\mathcal{A}\in\{0,1\}^{M\times N}$ represent the task assignment matrix, where $\mathcal{A}_{i,j}=1$ indicates that robot $i$ is assigned to track target $j$, and $\mathcal{A}_{i,j}=0$ otherwise. The vector $\mathbf{u}=(u_{1},\cdots,u_{M})$ denotes the control inputs or general actions for all robots. The original nonlinear mix-integer problem is nontrivial and cannot be updated at a higher frequency. To improve the efficiency and simplify the reasoning for the LLM, we decouple the problem as a bi-level optimization,
$\min_{\mathcal{A}}\quad$ $\mathcal{J}(\mathcal{A},\mathbf{u}^{*}),$ (2a)
s.t. $\mathbf{u}^{*}(\mathcal{A})\in\arg\min_{\mathbf{u}}\left\{\mathcal{J}(\mathcal{A},\mathbf{u})\mid\right.$ (2b)
$\qquad\qquad\mathcal{H}(\mathcal{A},\mathbf{u})=\bm{0},\ \mathcal{G}(\mathcal{A},\mathbf{u})\preceq\bm{0}\},$ (2c)
$\mathcal{H}(\mathcal{A},\mathbf{u}^{*})=\bm{0},$ (2d)
$\mathcal{G}(\mathcal{A},\mathbf{u}^{*})\preceq\bm{0}.$ (2e)
The high-level problem usually evaluates the task performance based on the cost of deployments towards robots, as discussed in [35]. The subproblem focuses on optimizing the robots’ control commands, given the assignment matrix generated by the high-level problem. Hence, the low-level optimization can be formulated as
$\min_{\mathbf{u},\bm{\nu},\bm{\xi}}$ $\ w_{1}f(\mathbf{x}_{t+1})+w_{2}\lVert\mathbf{u}\rVert+w_{3}\lVert\bm{\nu}\rVert+w_{4}\lVert\bm{\xi}\rVert,$ (3a)
s.t. $\quad\mathbf{x}_{t+1}=g(\mathbf{x}_{t},\mathbf{u}),\quad$ (3b)
$\quad\mathcal{G}(\mathbf{x}_{t+1},\mathcal{S})\preceq\bm{\nu},$ (3c)
$\quad\mathcal{G}(\mathbf{x}_{t+1},\mathcal{C})\preceq\bm{\xi},$ (3d)
$\quad\mathbf{u}\in\mathcal{U},\quad\mathbf{x}_{t+1}\in\mathcal{X}_{t+1}^{\text{free}},\quad\bm{\nu},\bm{\xi}\succeq\mathbf{0},$ (3e)
where the objective function is composed of four components weighted by $w_{1},w_{2},w_{3},w_{4}\in\mathbb{R}_{+}$. The equation (3b) represents the dynamic model of robots, and $\mathbf{x}_{t}$ is the robot state at time $t$. The constraint (3c) denotes the probabilistic safe constraints toward known sensing danger zones $\mathcal{S}=(\mathcal{S}_{1},\cdots,\mathcal{S}_{p})$, and (3d) represents the safety conditions for known communication danger zones $\mathcal{C}=(\mathcal{C}_{1},\cdots,\mathcal{C}_{q})$ with slack variables $\bm{\nu},\bm{\xi}$ for relaxation.
Instead of deploying every LLM for each individual agent, we adopt a partially centralized optimization, where task LLM guides task assignment, and action LLM assists in tracking processes. The design of the prompts and their corresponding verification methods will be discussed in the following sections.
III-C Task LLM for Reconfiguration
We query task LLM to output the assignment matrix to avoid optimization formulation with matrix-based hard constraints, which is not straightforward to reason. Task assignment in a multi-robot tracking problem requires comprehensive information about robot capabilities, available resources, and the current status of both robots and dynamic tasks (targets). Additionally, historical data is important to evaluate whether a robot’s assignment should be updated. The following information is provided to the task LLM in a descriptive language format: (1) robot and target status: This includes the dynamic states of robots (e.g., position, velocity), target states, and the current status of robots failures by attacks (e.g., whether they are affected by sensing or communication danger zones). (2) environmental status: Known locations of sensing and communication danger zones, along with their impact on the robots. (3) robot capacities: The number of targets a robot can track at once, which may also be dynamically adjusted based on current conditions. (4) historical data: The most recent $k$ results about states of robots and targets, information on previous danger zone attacks, and prior task assignments. This historical context helps the LLM make informed decisions about potential reassignment needs.
III-C1 Prompt Design
The system prompt is designed to feed into LLM by clearly defining its role as an optimization engine tasked with assigning tracking targets to robots. You are an optimizer with the goal of assigning robots to track targets. The user input provides essential real-time and historical information about robots, targets, and the environment. An example is provided:
The recent k results of status and observations are ... The []th information is as follows. Drones are currently at the following positions: Drone [ID] is at [] .... Targets are currently at the following positions: Target [ID] is at [ ] ... The known sensing zones are: Drone [ID] knows the sensor zone [ID] located at ... No attack has been detected. The last trace of the tracking estimation covariances matrix is []. The optimization constraints and request for LLM outputs are described in the form of
Each drone has the ability to track at most [] targets, and each target should be tracked by at least one drone as possible. Please provide the new tracking assignment for each drone in each line with ’’. This structured approach ensures that the LLM has all necessary data, constraints, and output expectations.
III-C2 Output Verification
We employ two types of checkers: the first ensures that the LLM output adheres to the required format with numerical data, while the second verifies whether the output meets the specified hard constraints. The verification function is designed to ensure that the output from the LLM can be used effectively for low-level optimization. It helps to determine whether the current output should be used or skipped, adjusting the prompt for the next query if needed. The assignment matrix is evaluated using with
$\beta(\mathcal{A})=\mathbf{1}^{T}\max\left(\mathbf{0},\mathcal{A}\mathbf{1}-\mathbf{c}\right)+\mathbf{1}^{T}\max\left(\mathbf{0},\mathbf{1}-\mathcal{A}^{T}\mathbf{1}\right)$ (4)
where $\mathbf{c}=(c_{1},\cdots,c_{M})$ is the vector of robot capacities, $\mathbf{1}$ represent a vector of ones. The first term evaluates the robot capacity constraint as satisfying, where $c_{i}$ is the maximum number of targets the robot $i$ can track. The second term qualifies as each target being tracked by at least one robot. If $\beta(\mathcal{A})>0$, then the constraints are not met and we will skip the current output.
III-D Action LLM for Performance Adaptation
III-D1 Prompt Design
We specify the reasoning meaning of different objectives to the action LLM to evaluate their importance and adjust weights $(w_{1},w_{2},w_{3},w_{4})$ accordingly. The description for each objective is:
1. tracking error computed by the trace of the estimation covariance matrix of the targets 2. control cost computed by the norm of control input 3. slack variables of safety constraints to avoid sensing danger zones 4. slack variables of safety constraints to avoid communication danger zones We use action LLM to evaluate the relative importance of each objective and adjust the parameters $w_{1},w_{2},w_{3},w_{4}$ accordingly. The weight adjustment is a short-horizon decision. Therefore, we provide the LLM with the current status of robots, targets, and environments, with the last weight values. Similarly, the system prompt is designed as:
You are a multiple objective optimizer with the goal of specifying the weights of each objective function. In addition to status, the weight is provided with the specified output format.
The current weights for objective functions are: ... You should give a new weight as a list with a length of 4.
III-D2 Output Verification
The weights in our optimization framework typically indirectly influence the results. To determine whether the output generated by the LLM is appropriate for optimization, we add empirical bounds for each weight. We use such bounds to filter the LLM’s results, ensuring that only outputs meeting predefined criteria are passed on for further optimization.
III-E Human Input
We consider human reactions to supervise the system to inspect abnormal behaviors, risks, and uncertainty. Humans are generally not able to directly manage massive numerical data in real-time, like the exact positions of all robots, targets, and danger zones, but can have a sense of the actions of robots given some visualizations of data and real-world unmodeled conditions. The human input is optional for the system, and only triggered and added to the user prompt if the human has comments about the current performance. The additional prompt is at the end of the scenario description, and starts with,
In addition, the human supervisor has some input... We list the following types of input (1) performance qualifications: Humans can provide questions or concerns about the general performance. For example, humans can specify one robot to conduct subtracting, like focusing on the track part of the targets. Humans can also raise questions about the performance, reporting the missing tracking of some target, zigzag actions of robots, etc. (2) risk assessment: For environmental hazards, humans can specify some high-level evaluation like to push the robot team to be more conservative, or detailed information like to avoid a particular danger region if the target is very close to this region. (3) abnormal specifies: When deploying the framework to hardware platforms, sometimes there are extra factors that might affect the performance. For example, a robot might get trapped in a local minimum and not move. Then, a human can specify that the robot doesn’t move for a time and what’s the reason for that. In addition, the danger zone we assume might be more severe than its model and thus needs priority adjustment for safety.
IV Results
IV-A Implement Details
We evaluated the capabilities of our framework through numerical simulations in different scenarios. For our experiments, we leveraged both the GPT-3.5-Turbo [36] and GPT-4o [37] models and configured the temperature to 0 to ensure deterministic responses. The maximum token limit is set to $50\times(\bar{c}+2)$ for the GPT-3.5-Turbo model, where $\bar{c}$ is the average capacity of the robots. For the GPT-4o model, we add an extra 200 tokens allowance as it usually generates more explanations in the output.
During the simulations, the inner LLM was updated every 2-3 steps to manage tuned performance, while the outer LLM was refreshed every 8-10 steps to handle higher-level task assignments. The simulation runs for 300 steps. In real-world experiments, the action LLM was updated every 5 seconds, while the task LLM was triggered every 20-30 seconds to assess task allocation and overall performance. The task LLM incorporated the status history from the past 5 steps to provide context for long horizon decision-making. For more scenarios and parameters, please refer to https://github.com/Zhourobotics/hierarchical-llms.git.
IV-B Simulations
IV-B1 Reasoning Ability
We first varied the number of targets and robots to determine whether LLMs can deliver reasonable solutions in the correct format as the task complexity increases. The tests are conducted in an environment with five sensing danger zones and two communication danger zones, using between two to eight robots and two to eight targets. Each robot’s tracking capability is fixed at half the number of targets, ranging from one to four, to ensure a balanced setup.
To study the performance of different LLM models, we carried out experiments with two different models, GPT-4o and GPT-3.5-Turbo. As a result, GPT-4o consistently produced feasible solutions with a 100% success rate, while the success rate of the GPT-3.5-Turbo model is illustrated in Fig. 4. Note that in all of our experiments, the full information of the system is given to the LLMs. This high success rate indicates the capability of LLMs in understanding and reasoning about constraints when global information is available. For the task LLM, we tested scenarios with and without providing examples of prior task assignments. Even in cases where no prior reference is given, the LLM still outputs feasible solutions that satisfy the necessary constraints, demonstrating its adaptability and reliability under increased task complexity. From the results of our experiments, the performance of Task LLM using the GPT-3.5-Turbo model deteriorates when a small number of robots have to track a large number of targets, which makes it more challenging to satisfy the requirements.
(a) Task LLM
(b) Action LLM
Figure 4: Success rate of both the task LLM (a) and the action LLM (b) with the number of robots and targets.
The average token number using these two models is shown in Fig. 5 and Fig. 6. As expected, the token count generally increases with the growing number of robots and targets due to the rising complexity of task assignments as more robots and targets are introduced. The task LLM, in particular, shows a noticeable increase in token usage with higher target numbers. The action LLM doesn’t have an obvious change because even if the prompt token increases, the output maintains a consistent level of information.
(a) Task LLM
(b) Action LLM
Figure 5: Average token count generated by the GPT-3.5-Turbo model for Task LLM (a) and Action LLM (b) with the number of robots and targets.
(a) Task LLM
(b) Action LLM
Figure 6: Average token count generated by the GPT-4o model for Task LLM (a) and Action LLM (b) with the number of robots and targets.
IV-B2 Modular Evaluation
We conducted a comparison of three approaches: one without any LLMs (w/o LLMs), another with LLMs in the loop (w/ LLMs), and one with both LLMs and human input into the system. We consider a scenario with two drones tracking four targets in the same direction. We specify targets to have different velocities to increase the complexity of the conditions. The environment includes two sensing danger zones and one communication danger zone. The human input is included at a lower frequency to push the LLMs to improve the performance, the prompt is
Focus more on tracking targets; The trace is not good. We utilized GPT-4o for better performance, with the results presented in Tab. I. By adding the LLMs in the loop to adjust the parameters, the system can achieve better tracking performance and get fewer attacks from sensing and communication danger zones. With human input to express concern about the performance, LLMs can dramatically decrease the tracking errors computed by the trace of the estimation covariance matrix.
TABLE I: Comparison of Overall Performance
Accumlated
Trace
Sensing
Attacks
Comm.
Attacks
Trajectory
Length (m)
w/o LLMs 120.08 31 11 17.53
w/ LLMs 116.86 23 10 17.69
w/ LLMs + Human 100.98 19 7 18.55
IV-C Hardware Experiments
We validated the ability of the proposed framework through experiments with a team of four robots in both indoor and outdoor scenarios, as shown in Fig. 1 and 7. The outdoor experiments include more challenges like noisy detection, estimation, and control under disturbance. In particular, we set up two Scarab ground robots [38] to serve as moving targets, with each carrying an AprilTag [39] on top. To effectively detect and track the moving targets, two Dragonfly 230 MAVs [40] were employed. The Dragonfly 230 MAV is equipped with a downward-facing VGA camera for state estimation and target detection. An Unscented Kalman Filter is running on board the MAV to fuse the 30Hz VIO odometry from the VOXL board with the IMU to get 150Hz pose information. We used an additional AprilTag to create a common reference frame for all MAVs. Once an aerial agent observes the tag, the agent estimates the relative transformation between the common reference frame and its odometry frame and registers the detected target in the common reference frame. The robot receives global waypoints from the ground station and transforms them into its odometry frame for trajectory generation and control. The state estimation, tag detection, trajectory planning, and controller run onboard the drones.
(a) Indoor
(b) Outdoor
Figure 7: Hardware experiment of multi-drone target tracking with various danger zones (red disks represent sensing danger zones, and blue disks represent communication danger zones).
We created a series of human inputs during the experiments to see how they would impact the systems and help the LLMs better understand real-world conditions. These inputs focus on both performance and environmental concerns, allowing us to test the system’s responsiveness and adaptability. For performance input, we provide some examples of human expression about tracking performance, overall strategy, and individual robot behavior.
The robots should be more aggressive. Why does the Drone 21 look get stuck? The Drone 21 should focus on tracking target 122.
To address unexpected changes and concerns regarding environmental conditions or hazards, the following examples are provided:
The sensing danger zone causes more severe damage than we assume. The target 122 is in the danger zone.
In addition, a case study is shown below. Initially, the output specified tracking both targets, as follows:
1. Drone 21 should track Target 122 and Target 124. 2. Drone 26 should track Target 122 and Target 124.
However, tracking both targets simultaneously can lead to ambiguous behavior, making it unclear whether the drones are effectively monitoring the targets. Consequently, the human supervisor raised a concern,
The human supervisor has inputs: The robots seem to not track the targets.
In response, the Task LLM adjusted the assignment, refocusing each drone on a single target. The updated plan was as follows:
Based on the information provided, here is the new tracking assignment for each drone: - Drone 21 will track Target 122
- Drone 26 will track Target 124
Feedback to the human supervisor: The new tracking assignment has been optimized based on the capabilities and constraints of the drones, ensuring that each target is being tracked by at least one drone. The assignments have been made considering the capabilities and sensor failures of the drones.
This interaction process demonstrates how the system uses human input to reconfigure task allocations, improving real-time decision-making and operational efficiency.
V Conclusion
In this work, we propose a hierarchical framework that combines the reasoning capabilities of LLMs with traditional optimization techniques for multi-robot target tracking. Our results demonstrate the capacity of this framework to handle unexpected conditions, showing improved performance in resilience, adaptability, and tracking accuracy, particularly in dynamic and hazardous environments. This highlights the potential of LLM-integrated robotics systems in a wide range of real-world applications.